# Accelerated Parallel Tempering via Neural Transports

### v2

@@ C.2 Further Details on Control Accelerated PT
At the limit $K\rightarrow\infty$, following [Berner et al., 2025, Lemma B.7], and by the controlled Crooks fluctuation theorem [Vaikuntanathan and Jarzynski, 2008, Vargas et al., 2024], we arrive at the generalised work functional
@@
Note that the discrete version discussed in the main text is not the only discretisation choice. Other options [Albergo and Vanden-Eijnden, 2024, Máté and Fleuret, 2023] may also be applied.
@@ C.3 Further Details on Diff-APT
For a given annealing schedule $10=s_{0}<\ldots<s_{N}=$ 1, we construct $P_{k}^{n-1},Q_{k-1}^{n}$ through the linear discretisation of the time-reversal SDE on the time interval $[s_{n-1},s_{n}]$ with step size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$.
@@
We parametrise an energy-based model as outlined in Phillips et al. [2024] but modified to ensure that $\pi^{\theta}_{0}(x)\propto N(x;0,\mathrm{I})$. For completeness, we specifically take $\log\pi^{\theta}_{s}(x)=\log g^{\theta}_{s}(x)-\frac{1}{2}||x||^{2}$ where
@@ D.1 Target Distributions
We take the 40-mode Gaussian mixture model (GMM-2) in 2 dimensions from Midgley et al. [2023] where to extend this distribution to higher dimensions $d$, we extend the means with zero padding to a vector in $\mathbb{R}^{d}$ and keep the covariances as the identity matrix but now within $\mathbb{R}^{d}$. This helps to disentangle the effect of multi-modality from the effect of dimensionality on performance as we essentially fix the structure of the modes across different values of $d$. Following previous work [Akhound-Sadegh et al., 2024, Phillips et al., 2024], we also scale the distribution GMM-$d$ by a factor of 40 to ensure the modes are contained within the range $[-1,1]^{d}$ for our experiments with APT, where we also use the same scaling for the PT baseline to ensure a fair comparison.
@@
We take the DW-4 target from Köhler et al. [2020] which describes the energy landscape for a toy system of 4 particles $\{x_{1},x_{2},x_{3},x_{4}\}$ and $x_{i}\in\mathbb{R}^{2}$ given by
@@
We take the ManyWell-32 target from Midgley et al. [2023] formed from concatenating 16 copies of the 2-dimensional distribution
@@ D.2 Network and Training Details
For GMM-$d$, we use 20 RealNVP layers where we employ a 2-layer MLP with 128 hidden units for the scale and translation functions [Dinh et al., 2017]. We initialize the flow at the identity transformation. We use the Adam optimizer with a learning rate of 1e-3, perform gradient clipping with norm $1$ and employ EMA with decay parameter $0.99$.
@@
For both GMM-$d$ and MW-32, we use a 4-layer MLP with 512 hidden units, and for DW-4, we use a 4-layer EGNN [Satorras et al., 2021] with 64 hidden units. Recall that in CMCD-APT, we use the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. Therefore, our network is conditional on the $\beta$. We optimise the MLP by Adam with a learning rate of 1e-3 and EGNN with a learning rate of 1e-4. We additionally use gradient clipping with norm 1 for stability.
@@
  • Tuning PT: We define $\{\beta_{n}\}_{n=1}^{N}$ uniformly from 0 to 1. Then, we run PT for 600 steps, remove the first 100 steps as burn-in, and take the last 500 steps to calculate the rejection rate between adjacent chains. We then tune the value of $\beta_{n}$ to ensure the rejection rate even, according to Syed et al. [2021]. We repeat this process 10 times to ensure $\{\beta_{n}\}_{n=1}^{N}$ is stable.
@@
  • Testing: We follow Algorithm 1, running CMCD-APT for 100K steps, and calculate the round trip.
@@
For GMM-$d$, we use a 3-layer MLP with 128 hidden units, for MW-32, we use a 3-layer MLP with 256 hidden units and for DW-4, we use a 3 layer EGNN with 128 hidden units. For all models, we use a learning rate of 5e-4, gradient clipping with norm 1 for stability and EMA with decay parameter of 0.99. For training, we follow the same pipeline as CMCD-APT, but we only retain samples at the target chain in order to optimise the standard score matching objective. At sampling time, we tune the annealing schedule by running Diff-APT for 1,100 steps, discarding the first 100 samples as burn-in and using the last 1,000 steps to calculate rejection rates to apply the schedule tuning algorithm from Syed et al. [2021]. We then repeat this 10 times where we initialise from the uniform schedule.
@@ D.5 Further Details on Free Energy Estimator
We take the ground truth free energy of DW-4 from estimating $\Delta F$ with the APT estimator using 100,000 samples from PT with 60 chains (after tuning). We take the ground truth free energy of MW-32 from Midgley et al. [2023] which calculates the normalising constant of a single copy of $\hat{\pi}$ numerically allowing for the trivial computation of the overall normalising constant.
@@ D.6 Further Details on Comparing APT with Neural Samplers
For Diffusion, we follow the same procedure but we tune our annealing schedule with the same schedule tuning algorithm from Syed et al. [2021] as mentioned above (this step is not required for CMCD as the annealing schedule is required to be fixed during training) before we collect samples by applying $P_{k}^{n-1}$.
@@
D.7 License
@@ E.1 Visualisation of ManyWell-32
In Figures 5 and 6, we visualise ManyWell-32 samples generated by 1,000 consecutive CMCD-APT and Diff-APT steps with $N=5,10,30$ and $K=1,5$ from ManyWell-32 via marginal projections over the first four dimensions. We also show 1,000 independent ground truth samples in Figure 7 for comparison. We choose to report only 1,000 steps to illustrate how fast our sampler mixes. As we can see, the mode weights become more accurate as we increase $N$ and $K$.
@@ E.2 Comparison of Acceleration Methods for ManyWell-32 and DW-4
Method Neural Call ()($\downarrow$)( ↓ ) $\hat{\Lambda}$ ()($\downarrow$)( ↓ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ )
@@
Method Neural Call ()($\downarrow$)( ↓ ) $\hat{\Lambda}$ ()($\downarrow$)( ↓ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ )

### v3

@@ C.2 Further Details on Control Accelerated PT
At the limit $K\rightarrow\infty$, following (Berner et al., 2025, Lemma B.7), and by the controlled Crooks fluctuation theorem (Vaikuntanathan & Jarzynski, 2008; Vargas et al., 2024), we arrive at the generalised work functional
@@
Note that the discrete version discussed in the main text is not the only discretisation choice. Other options (Albergo & Vanden-Eijnden, 2024; Máté & Fleuret, 2023) may also be applied.
@@ C.3 Further Details on Diff-APT
For a given annealing schedule $0=s_{0}<\ldots<s_{N}=1$, we construct $P_{k}^{n-1},Q_{k-1}^{n}$ through the linear discretisation of the time-reversal SDE on the time interval $s_{n-1},s_{n}]$ with step size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$.
@@
[…]
@@ D.1 Target Distributions
We take the 40-mode Gaussian mixture model (GMM-2) in 2 dimensions from Midgley et al. (2023) where to extend this distribution to higher dimensions $d$, we extend the means with zero padding to a vector in $\mathbb{R}^{d}$ and keep the covariances as the identity matrix but now within $\mathbb{R}^{d}$. This helps to disentangle the effect of multi-modality from the effect of dimensionality on performance as we essentially fix the structure of the modes across different values of $d$. Following previous work (Akhound-Sadegh et al., 2024; Phillips et al., 2024), we also scale the distribution GMM-$d$ by a factor of 40 to ensure the modes are contained within the range $1,1]^{d}$ for our experiments with APT, where we also use the same scaling for the PT baseline to ensure a fair comparison.
@@
We take the DW-4 target from Köhler et al. (2020) which describes the energy landscape for a toy system of 4 particles $\{x_{1},x_{2},x_{3},x_{4}\}$ and $x_{i}\in\mathbb{R}^{2}$ given by
@@
We take the ManyWell-32 target from Midgley et al. (2023) formed from concatenating 16 copies of the 2-dimensional distribution
@@
Alanine Dipeptide
This is a small molecule with 22 atoms, each of which has 3 dimensions. The target energy is defined with the amber14/protein.ff14SB forcefield in vacuum using the DMFF library in JAX (Wang et al., 2023).
@@ D.2 Network and Training Details
For GMM-$d$, we use 20 RealNVP layers where we employ a 2-layer MLP with 128 hidden units for the scale and translation functions (Dinh et al., 2017). We initialize the flow at the identity transformation. We use the Adam optimizer with a learning rate of 1e-3, perform gradient clipping with norm $1$ and employ EMA with decay parameter $0.99$.
@@
For both GMM-$d$ and MW-32, we use a 4-layer MLP with 512 hidden units, and for DW-4, we use a 4-layer EGNN (Satorras et al., 2021) with 64 hidden units. Recall that in CMCD-APT, we use the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. Therefore, our network is conditional on the $\beta$. We optimise the MLP by Adam with a learning rate of 1e-3 and EGNN with a learning rate of 1e-4. We additionally use gradient clipping with norm 1 for stability.
@@
  • Tuning PT: We define $\{\beta_{n}\}_{n=1}^{N}$ uniformly from 0 to 1. Then, we run PT for 600 steps, remove the first 100 steps as burn-in, and take the last 500 steps to calculate the rejection rate between adjacent chains. We then tune the value of $\beta_{n}$ to ensure the rejection rate even, according to Syed et al. (2021). We repeat this process 10 times to ensure $\{\beta_{n}\}_{n=1}^{N}$ is stable.
@@
  • Testing: We follow Algorithm˜1, running CMCD-APT for 100K steps, and calculate the round trip.
@@
For GMM-$d$, we use a 3-layer MLP with 128 hidden units, for MW-32, we use a 3-layer MLP with 256 hidden units and for DW-4, we use a 3 layer EGNN with 128 hidden units. For all models, we use a learning rate of 5e-4, gradient clipping with norm 1 for stability and EMA with decay parameter of 0.99. For training, we follow the same pipeline as CMCD-APT, but we only retain samples at the target chain in order to optimise the standard score matching objective. At sampling time, we tune the annealing schedule by running Diff-APT for 1,100 steps, discarding the first 100 samples as burn-in and using the last 1,000 steps to calculate rejection rates to apply the schedule tuning algorithm from Syed et al. (2021). We then repeat this 10 times where we initialise from the uniform schedule.
@@ D.5 Further Details on Free Energy Estimator
We take the ground truth free energy of DW-4 from estimating $\Delta F$ with the APT estimator using 100,000 samples from PT with 60 chains (after tuning). We take the ground truth free energy of MW-32 from Midgley et al. (2023) which calculates the normalising constant of a single copy of $\hat{\pi}$ numerically allowing for the trivial computation of the overall normalising constant.
@@ D.6 Further Details on Comparing APT with Neural Samplers
For Diffusion, we follow the same procedure but we tune our annealing schedule with the same schedule tuning algorithm from Syed et al. (2021) as mentioned above (this step is not required for CMCD as the annealing schedule is required to be fixed during training) before we collect samples by applying $P_{k}^{n-1}$.
@@
D.7 Further Details on Alanine Dipeptide
We run CMCD-APT using 5 chains and $K=1,2,5$. For the local move, we adopt HMC with a dynamic step size: when the acceptance rate is larger than 0.9, we increase the step size by 1.2; if the acceptance rate is smaller than 0.8, we divide the step size by 1.2. Other settings are the same as those of other targets.
D.8 License
@@ E.1 Visualisation of ManyWell-32
In Figures˜5 and 6, we visualise ManyWell-32 samples generated by 1,000 consecutive CMCD-APT and Diff-APT steps with $N=5,10,30$ and $K=1,5$ from ManyWell-32 via marginal projections over the first four dimensions. We also show 1,000 independent ground truth samples in Figure˜7 for comparison. We choose to report only 1,000 steps to illustrate how fast our sampler mixes. As we can see, the mode weights become more accurate as we increase $N$ and $K$.
@@ E.2 Comparison of Acceleration Methods for ManyWell-32 and DW-4
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$